# Coexistence of Topological Anderson Insulator and Multifractal Critical Phase in a Non-Hermitian Quasicrystal

### v1

@@ Abstract
The interplay of topology, disorder, and non-Hermiticity gives rise to phenomena beyond the conventional classification of quantum phases. We propose a one-dimensional non-Hermitian Su–Schrieffer–Heeger model with quasiperiodically modulated nonreciprocal intracell hopping. We show that quasiperiodic modulation can substantially enhance topological robustness and, remarkably, induce a non-Hermitian topological Anderson insulator (TAI) phase. Beyond the topological transition, increasing nonreciprocity drives a cascade of localization transitions in which all bulk eigenstates evolve from extended to multifractal critical and ultimately to localized states. Strikingly, the extended-to-critical transition coincides exactly with a real–complex spectral transition. We establish complete phase diagrams and derive exact analytical boundaries for both topological and localization transitions, uncovering an unanticipated coexistence of TAI and multifractal critical phases. Finally, we propose a feasible implementation in topolectrical circuits. Our results reveal a new paradigm for the cooperative effects of topology, quasiperiodicity, and non-Hermiticity.
@@
Introduction-The discovery of topological phases has fundamentally reshaped our understanding of quantum matter, revealing boundary states protected by global topological invariants [1, 2, 3]. While this framework was originally formulated for Hermitian systems, growing interest in open and dissipative platforms has propelled the development of non-Hermitian topology [4, 5, 6, 7]. Non-Hermitian Hamiltonians naturally emerge in systems with gain and loss [8, 9], finite quasiparticle lifetimes [10], complex refractive indices in photonics [11, 12], and engineered Laplacians in electrical circuits [13, 14, 15]. In contrast to Hermitian counterparts, such systems host intrinsically non-Hermitian topological structures, including exceptional points [16] and complex-energy braiding [17, 18, 19]. Remarkably, they also support phases without Hermitian analogs [6, 7]. A defining feature of nonreciprocal systems is the breakdown of conventional bulk–boundary correspondence due to the non-Hermitian skin effect (NHSE), where an extensive number of eigenstates accumulate at the boundaries [20, 21]. The NHSE itself has a topological origin, intimately connected to point-gap topology under periodic boundary conditions [22, 23].
@@
Disorder, ubiquitous in realistic systems, generally destabilizes topological phases through Anderson localization [24]. Yet disorder can also play a constructive role: sufficiently strong randomness may induce topological transitions from trivial phases, leading to the topological Anderson insulator (TAI) [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The interplay between disorder and non-Hermiticity further enriches localization physics, producing unconventional spectral and transport behavior [42, 43, 44, 45, 46, 47]. Extending the TAI paradigm to non-Hermitian systems has revealed a diverse landscape of disorder-driven topological transitions [48, 49, 50, 51, 52]. In parallel, quasiperiodic systems such as the André–Aubry–Harper model [53, 54] provide a deterministic route to localization transitions. With suitable quasiperiodic modulations or long-range hopping, quasicrystals can host mobility edges separating extended and localized states [55, 56, 57, 58, 59]. The associated critical states may proliferate into an intermediate multifractal phase, distinguished from both extended and localized regimes by its spectral statistics and wave-function scaling properties [60, 61, 62, 63, 64, 65, 66, 67]. Such critical phases are central to nonergodic dynamics, anomalous transport, and many-body localization phenomena [68, 69, 70, 71]. Topological states in quasiperiodic lattices have also been extensively investigated [72, 73, 74, 75, 76, 77, 78, 79, 80]. Despite these advances, a fundamental question remains unresolved: can a disorder-induced topological Anderson phase and a multifractal critical phase coexist within a unified non-Hermitian quasiperiodic framework? To date, these phenomena have largely been treated as distinct regimes, and the conditions enabling their simultaneous emergence remain unknown.
@@
In this work, we resolve this question by introducing a one-dimensional non-Hermitian Su–Schrieffer–Heeger (SSH) model with quasiperiodically modulated nonreciprocal intracell hopping. We show that quasiperiodicity not only enhances the robustness of the topological phase but also drives a disorder-induced transition into a non-Hermitian TAI phase. More strikingly, increasing nonreciprocity induces a cascade of localization transitions, with all bulk eigenstates evolving from extended to multifractal critical and ultimately to localized regimes. The extended-to-critical transition coincides exactly with a real–complex spectral transition, establishing a direct correspondence between spectral feature and wave-function criticality. By deriving exact analytical phase boundaries, we construct the complete phase diagram and uncover a previously unexplored regime where the TAI and multifractal critical phases coexist. Finally, we propose a feasible implementation in topolectrical circuits, providing an experimentally accessible platform to probe the interplay of non-Hermiticity, topology, and quasiperiodic criticality.
@@
Model Hamiltonian-We consider a one-dimensional SSH chain with quasiperiodically modulated nonreciprocal intracell hopping, described by
@@
The Hamiltonian takes a tridiagonal non-Hermitian form under open boundary conditions. For $|\lambda|<|v|$, it can be mapped to a Hermitian matrix via a similarity transformation $h=D^{-1}HD$ with diagonal $D$ [81](see Supplemental Material (SM) [82]). Consequently, $H$ is pseudo-Hermitian in the sense of Mostafazadeh [83], and its spectrum is entirely real in this regime. As illustrated in Fig. 1(a) and 1(b), the eigenenergies remain real for $|\lambda|<|v|$, while a real–complex spectral transition occurs at $|\lambda|=|v|$, beyond which the similarity transformation breaks down and complex eigenvalues emerge.
To characterize localization, we compute the inverse participation ratio (IPR) of the $n$th eigenstate, $\mathrm{IPR}_{n}=\sum_{j,\alpha}\left(|\psi_{n,j}|^{4}+|\psi_{n,j\alpha}|^{4}\right)$, with $\alpha=A,B$. In the thermodynamic limit, $\mathrm{IPR}_{n}\rightarrow 0$ for extended states and remains finite for localized states. We further extract the fractal dimension, $\mathrm{FD}_{n}=-\lim_{L\rightarrow\infty}\left[\ln\text{IPR}_{n}/\ln L\right]$, for which $\mathrm{FD}_{n}=1$ $(0)$ corresponds to extended (localized) states, while $0<\text{FD}<1$ signals multifractal criticality. Note that there is no NHSE in our model due to that the nonreciprocity is quasiperiodic [84].
Transfer matrix and Lyapunov exponent-To analyze localization induced by quasiperiodic nonreciprocity, we formulate the transfer-matrix approach and compute the Lyapunov exponent. From the stationary Schrödinger equation $H|\psi\rangle=E|\psi\rangle$, with eigenstate $|\psi\rangle=\sum_{j}a_{j}|j,A\rangle+b_{j}|j,B\rangle$, yields
@@
where $N$ is the number of unit cells and $\left\|.\right\|$ denotes the matrix norm. The localization length is given by $\xi=\gamma^{-1}(E)$; thus $\gamma(E)>0$ signals localization, while $\gamma(E)=0$ corresponds to extended or critical states.
Topological Anderson insulator-For the clean SSH model ($\lambda=0$), the system is topologically nontrivial when $|v|<|w|$. Introducing quasiperiodic nonreciprocity fundamentally alters this criterion. Remarkably, we find that a topological phase can emerge in the parameter regime that is trivial in the clean limit, thereby realizing a non-Hermitian topological Anderson insulator (TAI). Figure 1(c) shows the open-boundary spectrum $|E|$ as a function of $\lambda$ when $v=1.2>w$, where the clean system is topologically trivial. As $\lambda$ increases, zero-energy modes appear within a finite parameter window. Their spatial profiles [Fig. 1(f)] reveal two states exponentially localized at opposite ends of the chain, confirming the onset of a nontrivial topological phase induced by quasiperiodic nonreciprocity.
The topological transition can be determined analytically from the divergence of the localization length of the zero-energy modes. Setting $E=0$ in Eq. (3), the Lyapunov exponent reduces to (see details in SM [82])
$\gamma_{0}=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\ln\left|\frac{v-\lambda\cos\theta_{j}}{w}\right|.$ (8)
Using phase averaging based on Weyl’s equidistribution theorem and irrational rotations [85, 86, 87], we obtain
$\gamma_{0}=\left\{\begin{array}[]{ll}\ln\frac{\left|\lambda\right|}{2\left|w\right|},\qquad&\left|v\right|\leq\left|\lambda\right|;\\ \ln\frac{\left|v\right|+\sqrt{v^{2}-\lambda^{2}}}{2\left|w\right|},\qquad&\left|v\right|>\left|\lambda\right|.\end{array}\right$. (9)
@@
$|\lambda|=|2w|,\quad|v|+\sqrt{v^{2}-\lambda^{2}}=2|w|.$ (10)
These expressions provide exact analytical criteria for the disorder-induced topological transition.
To further characterize the topological phase, we compute the real-space winding number under open boundary conditions [88]. The Hamiltonian preserves chiral symmetry, $SHS^{-1}=-H$, with $S=\sigma_{z}\otimes\mathcal{I}$. Let $|\psi^{\pm}_{nR}\rangle$ and $|\psi^{\pm}_{nL}\rangle$ denote the biorthonormal right and left eigenstates satisfying $H^{\dagger}|\psi^{\pm}_{nL}\rangle=\pm E_{n}^{\star}|\psi^{\pm}_{nL}\rangle$ and $H^{\dagger}|\psi^{\pm}_{nL}\rangle=\pm E_{n}^{\star}|\psi^{\pm}_{nL}\rangle$. Excluding edge modes, we define the open-boundary $Q$ matrix as $Q=\sum_{n}\left(|\psi^{+}_{nR}\rangle\langle\psi^{+}_{nR}|-|\psi^{-}_{nL}\rangle\langle\psi^{-}_{nL}|\right)$, and the winding number
@@
where $X$ is the coordinate operator and $\mathrm{Tr^{\prime}}$ denotes the trace over a central region of length $L^{\prime}$. We find $W=1$ in the TAI phase and $W=0$ in the trivial regime.
The resulting phase diagram in the $v-\lambda$ plane is shown in Fig. 2(a). The numerically obtained winding number agrees precisely with the analytical boundaries in Eq. (10), confirming the emergence of a disorder-induced topological phase.
Figure 1: (Color online) (a) Real and (b) imaginary parts of the eigenenergies under periodic boundary conditions as a function of the nonreciprocity strength $\lambda$. The black and blue dashed lines mark the critical values $\lambda=v$ and $\lambda=2w$, separating the extended, multifractal critical, and localized regimes. The color scale indicates the fractal dimension $\mathrm{FD}$ of the corresponding eigenstates. (c) Energy spectrum $|E|$ under open boundary conditions. (d)–(e) Representative spatial profiles of eigenstates in the extended, critical, and localized phases, respectively. (f) Zero-energy edge modes in the topological Anderson insulating phase. Parameters: $v=1.2$, $w=1$, and system size $L=2N=1220$.
Figure 2: (Color online) (a) Phase diagram in the $v-\lambda$ plane showing the topological Anderson insulating (TAI) and trivial phases. The color scale represents the real-space winding number $W$. Red dashed and solid lines denote the analytical phase boundaries given by Eq. (10). (b) Phase diagram of the extended, multifractal critical, and localized regimes determined from the Lyapunov exponent and fractal dimension. The black dashed line indicates the cut at $v=1.2$.

### v2

@@ Abstract
The interplay of topology, disorder, and non-Hermiticity gives rise to phenomena beyond the conventional classification of quantum phases. We propose a one-dimensional non-Hermitian Su–Schrieffer–Heeger model with quasiperiodically modulated nonreciprocal intracell hopping. We show that quasiperiodic modulation can substantially enhance the topological regime and, remarkably, induce a non-Hermitian topological Anderson insulator (TAI) phase. Beyond the topological transition, increasing nonreciprocity drives a cascade of localization transitions in which all bulk eigenstates evolve from extended to multifractal critical and ultimately to localized states. Strikingly, the extended-to-critical transition coincides exactly with a real–complex spectral transition. We establish complete phase diagrams and derive exact analytical boundaries for both topological and localization transitions, uncovering an unanticipated coexistence of TAI and multifractal critical phases. Finally, we propose a feasible implementation in topolectrical circuits. Our results reveal a new paradigm for studying the cooperative effects of topology, quasiperiodicity, and non-Hermiticity.
@@ I Introduction
The discovery of topological phases has fundamentally reshaped our understanding of quantum matter, revealing boundary states protected by global topological invariants [1, 2, 3]. While this framework was originally formulated for Hermitian systems, growing interest in open and dissipative platforms has propelled the development of non-Hermitian topology [4, 5, 6, 7]. Non-Hermitian Hamiltonians naturally emerge in systems with gain and loss [8, 9], finite quasiparticle lifetimes [10], complex refractive indices in photonics [11, 12], and engineered Laplacians in electrical circuits [13, 14, 15]. In contrast to Hermitian counterparts, such systems host intrinsically non-Hermitian topological structures, including exceptional points [16] and complex-energy braiding [17, 18, 19]. Remarkably, they also support phases without Hermitian analogs [6, 7]. A defining feature of nonreciprocal systems is the breakdown of conventional bulk–boundary correspondence due to the non-Hermitian skin effect (NHSE), where an extensive number of eigenstates accumulate at the boundaries [20, 21]. The NHSE itself has a topological origin, intimately connected to point-gap topology under periodic boundary conditions [22, 23, 24].
@@
Disorder, ubiquitous in realistic systems, generally destabilizes topological phases through Anderson localization [25]. Yet disorder can also play a constructive role: sufficiently strong randomness may induce topological transitions from trivial phases, leading to the topological Anderson insulator (TAI) [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. The TAI phases and the topological properties based on the 1D Su–Schrieffer–Heeger (SSH) model with quasiperiodic disorders have also been studied extensively [43, 44, 45, 46, 47, 48]. The interplay between disorder and non-Hermiticity further enriches localization and topological physics, producing unconventional phenomena that cannot be observed in the Hermitian systems [49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. Extending the TAI paradigm to non-Hermitian systems has revealed a diverse landscape of disorder-driven topological transitions [61, 62, 63, 45, 64]. In parallel, quasiperiodic systems such as the André–Aubry–Harper model [65, 66] provide a deterministic route to localization transitions. With suitable quasiperiodic modulations or long-range hopping, quasicrystals can host mobility edges separating extended and localized states [67, 68, 69, 70, 71]. The associated critical states may proliferate into an intermediate multifractal phase, distinguished from both extended and localized regimes by its spectral statistics and wave-function scaling properties [72, 73, 74, 75, 76, 77, 78, 79]. Such critical phases are central to nonergodic dynamics, anomalous transport, and many-body localization phenomena [80, 81, 82, 83]. Topological states in such quasiperiodic lattices have also been extensively investigated [84, 85, 86, 87, 88, 89, 90, 91, 92]. Despite these advances, a fundamental question remains unresolved: can a disorder-induced TAI phase and a multifractal critical phase coexist within a unified non-Hermitian quasiperiodic framework? To date, these phenomena have largely been treated as distinct regimes, and the conditions enabling their simultaneous emergence remain unknown.
@@
In this work, we resolve this question by introducing a one-dimensional non-Hermitian SSH model with quasiperiodically modulated nonreciprocal intracell hopping. We show that quasiperiodicity not only enhances the robustness of the topological phase but also drives a disorder-induced transition into a non-Hermitian TAI phase. More strikingly, increasing nonreciprocity induces a cascade of localization transitions, with all bulk eigenstates evolving from extended to multifractal critical and ultimately to localized regimes. The extended-to-critical transition coincides exactly with a real–complex spectral transition, establishing a direct correspondence between spectral feature and wave-function criticality. By deriving exact analytical phase boundaries, we construct the complete phase diagram and uncover a previously unexplored regime where the TAI and multifractal critical phases coexist. Finally, we propose a feasible implementation of our model in topolectrical circuits. Our work reveals the coexistence of a non-Hermitian TAI phase and a multifractal critical phase for all bulk states, where both the topological and the extended-critical-localized phase boundaries are analytically tractable. Hence our model provides an important platform to study the interplay of non-Hermiticity, topology, and quasiperiodic disorders.
The rest of the paper is organized as follows. In Sec. II, we will first introduce the model Hamiltonian of the 1D non-Hermitian SSH model with quasiperiodically modulated nonreciprocal intracell hopping. In Sec. III, we discuss the topological Anderson insulating phase and determine the phase boundaries. Then we will further investigate the extended-critical-localized phase transitions in Sec. IV. We also propose an experimental scheme for realizing the model in electric circuit in Sec. V. Finally in Sec. VI, we will summarize our results.
II Model Hamiltonian
@@
We consider a one-dimensional SSH chain with quasiperiodically modulated nonreciprocal intracell hopping, described by
@@
The Hamiltonian takes a tridiagonal non-Hermitian form under open boundary conditions. For $|\lambda|<|v|$, it can be mapped to a Hermitian matrix via a similarity transformation $h=D^{-1}HD$ with diagonal $D$ [93](also see Appendix for details). Consequently, $H$ is pseudo-Hermitian in the sense of Mostafazadeh [94], and its spectrum is entirely real in this regime. Due to the quasiperiodicity of the nonreciprocal hopping, there will be no skin effect in this non-Hermitian model, and the spectra under open and periodic boundary conditions are consistent with each other [95]. As illustrated in Fig. 1(a) and 1(b), the eigenenergies remain real for $|\lambda|<|v|$ under periodic boundary conditions, while a real–complex spectral transition occurs at $|\lambda|=|v|$, beyond which the similarity transformation breaks down and complex eigenvalues emerge.
To analyze the localization induced by quasiperiodic nonreciprocity, we formulate the transfer-matrix approach and compute the Lyapunov exponent. From the stationary Schrödinger equation $H|\psi\rangle=E|\psi\rangle$, with eigenstate $|\psi\rangle=\sum_{j}a_{j}|j,A\rangle+b_{j}|j,B\rangle$, yields
@@
where $N$ is the number of unit cells and $\left\|.\right\|$ denotes the matrix norm. The localization length is given by $\xi=\gamma^{-1}(E)$; thus $\gamma(E)>0$ signals localized state, while $\gamma(E)=0$ corresponds to extended or critical state.
To numerically characterize the localization features of the eigenstates, we can also compute the moments $I_{q}(n)=\sum_{j,\alpha}|\psi_{nR,j\alpha}|^{2q}\propto L^{-D_{q}(q-1)}$ with $\alpha=A$ and $B$, where $\psi_{nR,j\alpha}$ is the $j$th component on sublattice $\alpha$ of the $n$th right eigenstate $|\psi_{nR}\rangle$ and $D_{q}$ are the fractal dimensions. $q$ is a positive integer. If the eigenstates are localized (extended), we have $D_{q}\rightarrow 0$ $(1)$. While for the multifractal states, we have $0<D_{q}<1$ [96, 70]. Taking $q=2$ in $I_{q}$, we obtain the famous inverse participation ratio (IPR) as $\mathrm{IPR}(\psi_{nR})=\sum_{j,\alpha}|\psi_{nR,j\alpha}|^{4}$ and fractal dimension as $D_{2}(\psi_{nR})=-\lim_{L\rightarrow\infty}\left[\ln\text{IPR}(\psi_{nR})/\ln L\right]$, which are extensively used to characterize the localization properties. In the thermodynamic limit, $\mathrm{IPR}(\psi_{nR})\rightarrow 0$ for extended states and remains finite for localized states.
Figure 1: (Color online) (a) Real and (b) imaginary parts of the eigenenergies under periodic boundary conditions as a function of the nonreciprocity strength $\lambda$. The black and blue dashed lines mark the critical values $\lambda=v$ and $\lambda=2w$, separating the extended, multifractal critical, and localized regimes. The color scale indicates the fractal dimension of the corresponding eigenstates. (c) Energy spectrum $|E|$ under open boundary conditions. (d)–(e) Representative spatial profiles of eigenstates in the extended, critical, and localized phases, respectively. (f) Zero-energy edge modes in the topological Anderson insulating phase. Parameters: $v=1.2$, $w=1$, and system size $L=2N=1220$.
III Topological Anderson insulator
For the clean SSH model ($\lambda=0$), the system is topologically nontrivial when $|v|<|w|$. Introducing quasiperiodic nonreciprocity fundamentally alters this criterion. Remarkably, we find that a topological phase can emerge in the parameter regime that is trivial in the clean limit, thereby realizing a non-Hermitian topological Anderson insulator (TAI). Figure 1(c) shows the open-boundary spectrum $|E|$ as a function of $\lambda$ when $v=1.2>w$, where the clean system is topologically trivial. As $\lambda$ increases, zero-energy modes appear within a finite parameter window. Their spatial profiles [Fig. 1(f)] reveal two states exponentially localized at opposite ends of the chain, confirming the onset of a nontrivial topological phase induced by quasiperiodic nonreciprocity.
The topological transition can be determined analytically from the divergence of the localization length of the zero-energy modes. Setting $E=0$ in Eq. (3), the two equations decouple and yield
$a_{j+1}=-\frac{v-\lambda\cos\theta_{j}}{w}a_{j}.$ (8)
Iterating from $j=1$ to $j=N$ gives
$a_{N+1}=(-1)^{N}\prod_{j=1}^{N}\frac{v-\lambda\cos\theta_{j}}{w}a_{1}.$ (9)
The Lyapunov exponent for the zero-energy mode is therefore
$\gamma(0)=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\ln\left|\frac{v-\lambda\cos\theta_{j}}{w}\right|.$ (10)
Since $\theta_{j}=2\pi\alpha j+\phi$ with irrational $\alpha$, Weyl’s equidistribution theorem ensures that the sequence is uniformly distributed on $0,2\pi)$, The Lyapunov exponent reduces to the phase average [97, 98, 53]
$\gamma(0)=\frac{1}{2\pi}\int_{0}^{2\pi}\ln\left|\frac{v-\lambda\cos\theta}{w}\right|d\theta.$ (11)
Using Jensen’s formula, the integral can be evaluated exactly, yielding
$\gamma(0)=\left\{\begin{array}[]{ll}\ln\frac{\left|v\right|+\sqrt{v^{2}-\lambda^{2}}}{2\left|w\right|},\qquad&\left|\lambda\right|<\left|v\right|;\\ \ln\frac{\left|\lambda\right|}{2\left|w\right|},\qquad&\left|\lambda\right|\geq\left|v\right|.\end{array}\right$. (12)
@@ III Topological Anderson insulator
$\begin{array}[]{ll}|v|+\sqrt{v^{2}-\lambda^{2}}=2|w|\qquad&\left|\lambda\right|<\left|v\right|;\\ |\lambda|=|2w|,\qquad&\left|\lambda\right|\geq\left|v\right|.\end{array}$ (13)
These expressions provide exact analytical criteria for the disorder-induced topological phase transition.
Figure 2: (Color online) (a) Phase diagram in the $v-\lambda$ plane showing the topologically nontrivial $(W=1)$ and trivial $(W=0)$ phases. The color scale represents the real-space winding number $W$. Red dashed and solid lines denote the analytical phase boundaries given by Eq. (13). (b) Phase diagram of the extended, multifractal critical, and localized regimes determined from the Lyapunov exponent and fractal dimension. The red dashed line is the same as in (a). The black solid line is the phase boundary separating the extended and localized phase. The black dashed line indicates the cut at $v=1.2$.
To further characterize the topological phase, we compute the real-space winding number under open boundary conditions [99]. The Hamiltonian preserves chiral symmetry, $SHS^{-1}=-H$, with $S=\sigma_{z}\otimes\mathcal{I}$. Let $|\psi^{\pm}_{nR}\rangle$ and $|\psi^{\pm}_{nL}\rangle$ denote the biorthonormal right and left eigenstates, which respectively satisfy the Schrödinger equations $H|\psi^{\pm}_{nR}\rangle=\pm E_{n}|\psi^{\pm}_{nR}\rangle$ and $H^{\dagger}|\psi^{\pm}_{nL}\rangle=\pm E_{n}^{\star}|\psi^{\pm}_{nL}\rangle$. Excluding edge modes, we define the open-boundary $Q$ matrix as $Q=\sum_{n}\left(|\psi^{+}_{nR}\rangle\langle\psi^{+}_{nL}|-|\psi^{-}_{nR}\rangle\langle\psi^{-}_{nL}|\right)$, and the winding number
@@
where $X$ is the coordinate operator and $\mathrm{Tr^{\prime}}$ denotes the trace over a central region of length $L^{\prime}$. In this work, we have taken the central region with $N/2\leq n\leq 3N/2$ such that $L^{\prime}=N$ in the numeric implementation. We find $W=1$ in the topologically nontrivial regime and $W=0$ in the trivial regime.
The resulting phase diagram in the $v-\lambda$ plane is shown in Fig. 2(a). We can see that in the region with $1<v<2$, the system is trivial when the disorder is weak. As $\lambda$ increases, the system will enter into the nontrivial phase with $W=1$, this is the topological Anderson insulating phase induced by the nonreciprocal quasiperiodic disorder. When $\lambda$ further increases, the system will become trivial again at $\lambda=2$. The boundaries separating the blue and yellow regions with different winding numbers, which are numerically obtained, agree precisely with the analytical boundaries in Eq. (13), as indicated by the red dashed and solid lines in Fig. 2(a), confirming the emergence of a disorder-induced topological phase.
IV Extended-critical-localized phase transitions